OpenStudy (anonymous):

Solve (y-6)^3+80=0 where y is a real number. Write your answer in simplified radical form. (If there is more than one solution, separate them with commas)

OpenStudy (anonymous):

\[(y-6)^3=-80\] take the cubed root of both sides \[y-6=\sqrt[3]{-80}\] then add 6 to both sides \[y=6+2i \sqrt[3]{10}, y=6-2i \sqrt[3]{10}\]] i believe this is what your looking for